21. The table to | Type of | Mass | Breathing Rate | Volume | % the right gives Animal (gm) (breaths/min)p respiratory parameters for 2 different species: one aquatic and A one terrestrial. Assume that atmospheric oxygen concentration is 02 used Oused per Extraction (mL/hr) breath (mL) 100 25 10 25 100 10 10 95 21%, and that water contains 34 ml O. at the conditions tested, what is the tidal volumes for species A & B? a) 9.5 ml b) 10 ml c) 95 ml Show steps please.
Answer #1

Tidal volume - body’s weight in kg/ respiratory rate* extraction *voulme per breath

Volume per breath of animal A = 25*10 =250ml

100gm=0.1kg

Tidal volume of animal A =0.1/25*25*250

10ml

Volume per breath of animal B = 10*10 =100ml

100gm=0.1kg

Tidal volume of animal B= 0.1/10*95*100

=95ml